OpenStudy (anonymous):

derivative: (x^3+1)^x (x^2+1)^(sinx)/x^(2cosx)

OpenStudy (anonymous):

I really hope you typed that wrong.

OpenStudy (anonymous):

nope, unfortunately that is the question :(

OpenStudy (amistre64):

id up that last one with a negative exponent and product rule it into submission

OpenStudy (amistre64):

\[D(fgh)=f'gh+fg'h+fgh'\]

OpenStudy (anonymous):

thank you, i'll try that now

OpenStudy (amistre64):

i think it makes life simpler in that you only have to focus on one part at a time ....

OpenStudy (anonymous):

or try logarithmic differentiation...

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

but that will be just as ugly

OpenStudy (amistre64):

you still gotsta play the logs in the product; but no matter which route you go, just know that someone hates you enough to make you do this kind of problem :)

OpenStudy (anonymous):

i've been working at it for 30 mins and it's so long and i got lost

OpenStudy (anonymous):

and confused

OpenStudy (anonymous):

Is this calc III?

OpenStudy (anonymous):

cal 1

OpenStudy (amistre64):

single variable derivatives are calc1

OpenStudy (anonymous):

Yea but it is a hell of a problem, we never had anything like this is my calc 1 class.

OpenStudy (amistre64):

georgia math, gotta love it :)

OpenStudy (anonymous):

i know! I gave up halfway. it's so long and confusing. and frustrating.

OpenStudy (anonymous):

you're from GA?

OpenStudy (amistre64):

florida, but a had a foreman from georgia and i used to make fun of him for the way he would do his math. it was some backwards backwoods sorta stuff but he always got the right answer :)

OpenStudy (anonymous):

haha

OpenStudy (amistre64):

y = (x^3+1)^x ln(y = (x^3+1)^x) ln(y) = x ln(x^3+1) y'/y = ln(x^3+1) + 3x^3/(x^3+1) ------------------- y = (x^2+1)^(sinx) ln(y=(x^2+1)^(sinx)) ln(y) = ln(x^2+1) sin(x) y'/y = ln(x^2+1) cos(x) + 2x sin(x)/(x^2+1) ---------------------- y = x^(-2cosx) ln(y = x^(-2cosx)) ln(y) = -2cos(x) ln(x) y'/y = 2sin(x) ln(x) -2cos(x)/x

OpenStudy (amistre64):

im not sure how far id try to simplify it after i put all the parts back in tho

OpenStudy (anonymous):

i'll just skip this question. haha. thank you though